OpenStudy (anonymous):

6(4-5)-2(4x-19)=15x+15 X=?

OpenStudy (anonymous):

multiply out the parenthesis. -6 - 8x +38 = 15x +15 add 8x to each side 32 = 23x +15 subtract 15 from each side 23x = 17 divide by 23

OpenStudy (anonymous):

x = 17/23 or .74 as a rounded decimal